OpenStudy (anonymous):

Solve for n. 7n = 35 A. 0.2 B. 5 C. 28 D. 42

OpenStudy (jango_in_dtown):

divide both sides by 7 and you have n=5

OpenStudy (alexandervonhumboldt2):

divide both sides by 7: 7n/7=35/7 n=5

OpenStudy (chantysquirrel1129**):

The first step you would do it divide both sides by 7 c:

OpenStudy (chantysquirrel1129**):

dang it wow xD

rebeccaxhawaii (rebeccaxhawaii):

b

OpenStudy (thread252):

n=5

OpenStudy (alexandervonhumboldt2):

direct answers!!

TheSmartOne (thesmartone):

._.

OpenStudy (chantysquirrel1129**):

You gave a direct answer, alex cx

OpenStudy (anonymous):

is it 5? b

TheSmartOne (thesmartone):

The easiest thing you could have done would have been to plug in your answer choices

OpenStudy (alexandervonhumboldt2):

lol

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (chantysquirrel1129**):

yes c:

OpenStudy (chantysquirrel1129**):

Alex, practice what you preach boi cx

rebeccaxhawaii (rebeccaxhawaii):

ITS B

OpenStudy (chantysquirrel1129**):

TSO, that's rude!

rebeccaxhawaii (rebeccaxhawaii):

ITS B

rebeccaxhawaii (rebeccaxhawaii):

ITS B

OpenStudy (anonymous):

Which equation can be used to solve the problem? How many 12-packs of juice boxes contain a total of 84 juice boxes? A. 12 + b = 84 B. 12 ÷ b = 84 C. 12b = 84 D. b ÷ 12 = 84 what baout this one?

OpenStudy (thread252):

yup the answer is b

OpenStudy (anonymous):

Which equation can be used to solve the problem? How many 12-packs of juice boxes contain a total of 84 juice boxes? A. 12 + b = 84 B. 12 ÷ b = 84 C. 12b = 84 D. b ÷ 12 = 84

OpenStudy (thread252):

C

OpenStudy (anonymous):

whats that one

OpenStudy (thread252):

answer is C

OpenStudy (chantysquirrel1129**):

She was probably typing it while everyone was like a a aaf So be nice >.>

TheSmartOne (thesmartone):

One question per post!

OpenStudy (anonymous):

Which equation can be used to solve the problem? How many 12-packs of juice boxes contain a total of 84 juice boxes? A. 12 + b = 84 B. 12 ÷ b = 84 C. 12b = 84 D. b ÷ 12 = 84

OpenStudy (anonymous):

???

OpenStudy (chantysquirrel1129**):

Post it on another post, emie c:

TheSmartOne (thesmartone):

\(\bf\Huge One~ question~ per~ post\)

OpenStudy (rccb):

C

OpenStudy (chantysquirrel1129**):

Lol you can't get any more clear than that xD